# Computational Obfuscations and Random Oracles for Derandomizing Asynchronous Consensus

### v1

@@ 6 Computational obfuscation for threshold encoding
2       $C[1,...,\ell][1,...m]$ an uninitizalized array of bit strings
@@ 5             $P[i]\leftarrow$ random bit string of length $k$
6             $r_{i}\leftarrow$ random bit string of length $m-k$
@@
8            if $i≥T$ then
9                   $a_{i}\leftarrow 1$
10
11            else
12                   $a_{i}\leftarrow 0$
13
             // enforce that XOR of $r_{i}$ is equal to $f_{t}(i)$
14             if $\bigoplus_{j=1}^{m-k}[r_{i}]_{j}\neq a_{i}$ then
15                   flip a bit in $r_{i}$ at random
16
17            $y_{i}\leftarrow P[i]\|r_{i}$
18             $C[i]\leftarrow H_{m}(y_{i})$
19
20      return $C,P$
21
Algorithm 6.1 Threshold Encapsulation
Preprocessing Algorithm.
Algorithm 6.1 is executed by the programmer to obtain a computational obfuscated protocol using preprocessing. The integer parameters $\ell$, $T$, $k$, and $m$ are declared (Line 1). Where $\ell$ is the span in which a threshold $t$ is defined, $m$ is the number of bits used by the one-way function, say 256 for SHA 256, and $k$ is the length of the (exposed) prefix nonce of the $m$ bits input (and output) of the SHA.
The for loop (lines 6.1 to 6.1) iterates over all possible values of $i$ and computes the cryptographic hash (e.g., the common one-way function such as SHA) over the nonce concatenated with $r_{i}$, a concealed random string of length $m-k$. The computation starts (line 6.1) with a random selection of bits string of size $k$ as the nonce for $i$ and entry and stores it in $P[i]$. Next, the algorithm randomly selects a bit string $r_{i}$ (line 6.1); this string must have the property that $\oplus_{j}[r_{i}]_{j}=1$ if and only if $i≥T$, and thus the algorithm sets $a_{i}=f(i)$ accordingly (lines 6.1 to 6.1), and checks (line 6.1) if the randomly chosen $r_{i}$ satisfies this property. If it does not, it simply flips one bit at random in $r_{i}$ (line 6.1). Since $\oplus_{j}[r_{i}]_{j}\neq a_{i}$, flipping any bit $[r_{i}]_{l}\leftarrow\neq[r_{i}]_{l}$ will cause the xor of the bits in $r_{i}$ to change and thus we get $\oplus_{j}[r_{i}]_{j}=a_{i}$.
The algorithm concatenates the nonce and $r_{i}$ to obtain a bit string $y_{i}$ of length $m$ (line 6.1), then the algorithm computes the hash of $y_{i}$ using $H_{m}$ and stores the resulting hash in $C[i]$ (line 6.1). By the time the algorithm reaches the return command (line 6.1), it has computed the hashes for all the $n$ hash values of $C$, and all the nonces used to calculate them are stored in $P$; thus, the algorithm returns them to be embedded in the protocol program.
local data: integers $\ell,m,k$, precomputed array of hashes $C[1,...,\ell][1,...,m]$ and nonces $P[1,...,\ell][1,...,k]$
[…]
2       $A[1,...,m]$ an uninitialized bit string
3       $A\leftarrow$ pre-image of hash $C[i]$ of size $m$ s.t., $A[1,...,k]=P[i][1,...,k]$
4 return $\bigoplus_{j=k}^{m}A[j]$
@@ 18             $C[i]\leftarrow H_{m}(y_{i})$
Execution of the Obfuscated Program.
The processes use the probingThreshold procedure described in Algorithm 6.2 to compute $f_{T}(i)$.
@@
The algorithm works by finding the pre-image $P[i]\|r_{i}$ that was used to compute $C[i]$ in the preprocessing phase. It begins by initializing a bit string $A$ of length $m$ to hold potential pre-images to $C[i]$ (line 2). Next, it finds and stores in $A$ a pre-image for $C[i]$ such that the first $k$ bits of the pre-image $A$ are equal to $P[i]$ (line 6.2) since if $A=P[i]\|r_{i}$ then $A[1,...,k]=P[i][1,...,k]$. This means that, with high probability $A=P[i]\|r_{i}$ and therefore $A[k,...,m]=r_{i}$. The exact method of finding this pre-image has not been specified, but exhaustive search mining or any other method to find a pre-image of a cryptographic hash function can be employed. Thus, Algorithm 6.2 returns $a_{i}$ by computing $\bigoplus_{j=k}^{m}A[j]$ (line 6.2), since the preprocessing algorithm defined $a_{i}=f_{T}(i)$, the algorithm returns $f_{T}(i)$.
Security Arguments Sketch.
Consider series of hashes $C=C_{1},\dots,C_{\ell}$ and $P=P_{1},\dots,P_{\ell}$ and nonces generated by Algorithm 6.1 for some $\ell,T,m,k$. We call the tuple $\mathcal{S}=(\ell,m,k,C,P)$ an obfuscation of the function $f_{T}$. We define a query to $\mathcal{S}$ with input $i\in[n]$ as finding the pre-image $P[i]\|r_{i}$ of a hash $C_{i}\in C$ and computing $\bigoplus_{j}[r_{i}]_{j}$.
Consider a malicious computationally bounded adversary that implements an adaptive adversary scheduler commonly assumed in the scope of asynchronous distributed computing. The scheduler computation bound is used to prove the security of the obfuscated construction. If the adversary can reveal $T$ in an expected number of $o(\log\ell)$ pre-images computations, then one can implement a search better than the binary search, which is proven optimal, a contradiction that implies the need to compute expected $\Theta(\log\ell)$ pre-images.
Furthermore, if the adversary can find a pre-image earlier than planned by (the SHA based time lock puzzle), then the one-way function used, say, SHA, is easy to reverse, which violates the common standard and belief.
7 From conciliators to consensus
In Section 5, we gave an implementation of a deterministic conciliator-like protocol where each process carries out $O(s)$ steps and a computation with $O(t)$ time complexity, while the adversary is forced to spend $Ω(st)$ time to find a bad execution that prevents agreement. This implementation requires significant setup, and because each instance is fixed and deterministic, once the adversary finds a bad execution, it costs nothing to reuse this bad execution if we are foolish enough to reuse the conciliator. We are also faced with the issue that we really want a protocol that guarantees agreement in all executions.
We can address both of these issues by supplementing the conciliator with a chain of adopt-commit and conciliator objects as suggested in [3]. An adopt-commit object [17, 2, 4] implements the adopt-commit protocol of Gafni [11], in which each process arrives with an input value $v$ and leaves with an output that is either $\left\langle{\textnormal{{adopt}},v^{\prime}}\right\rangle$ or $\left\langle{\textnormal{{commit}},v^{\prime}}\right\rangle$ for some value $v^{\prime}$. It satisfies the same validity and wait-freedom conditions as consensus, but agreement is replaced with two new conditions: consistency, which requires that if any process returns $\left\langle{\textnormal{{commit}},v}\right\rangle$, all other processes return $\left\langle{\textnormal{{commit}},v}\right\rangle$ or $\left\langle{\textnormal{{adopt}},v}\right\rangle$ for the same value $v$; and coherence, which requires that if all processes have the same input $v$, they all return $\left\langle{\textnormal{{commit}},v}\right\rangle$.
When used in alternation with conciliators, adopt-commit objects allow processes to decide immediately on any value that they see in a $\left\langle{\textnormal{{commit}},v}\right\rangle$, since consistency implies that all other processes will at least adopt the same value and the validity of the conciliators will preserve this agreement until the next adopt-commit, which will cause any remaining processes to also return $v$ because of coherence. So the only issue left is how to build conciliators that will push the processes into agreement.
For the first round, we use the threshold conciliator from Section 5. If this fails to produce agreement, in subsequent rounds we switch to a simpler random-oracle-based conciliator, in which a computationally-expensive hash function $H$ is applied to a nonce and the current round number to generate a value that acts like a shared coin agreed on by all the processes.
1procedure $\textnormal{{consensus}}(\textnormal{{input}},f,s,\textnormal{{nonce}})$
       // Attempt to reach agreement with thresholdConciliator
2       $v\textleftarrow\textnormal{{thresholdConciliator}}(\textnormal{{input}},f,s)$
3       $(a,v)\textleftarrow\textnormal{{adoptCommit}}_{1}(v)$
4       if $a=\textnormal{{commit}}$ then
5             return $v$
6
      // Switch to multi-round backup protocol
7       for $r\textleftarrow 2\dots∞$ do
             // Attempt to reach agreement with oracleConciliator
8             $v\textleftarrow\textnormal{{oracleConciliator}}(r,v,\textnormal{{nonce}})$
9             $(a,v)\textleftarrow\textnormal{{adoptCommit}}_{r}(v)$
10             if $a=\textnormal{{commit}}$ then
11                   return $v$
12
13
14
[…]
16       $\textnormal{{present}}[r][v]\textleftarrow 1$
17       if $\textnormal{{present}}[r][¬v]=0$ then
             // no disagreement, keep current preference
18             return $v$
19      else
             // disagreement, use random oracle
20             return $H(r,\textnormal{{nonce}})$
21
Algorithm 7.1 Derandomized consensus protocol
Pseudocode for this procedure is given in Algorithm 7.1. The main procedure consensus assumes an infinite array of shared adopt-commit objects $\textnormal{{adoptCommit}}_{1},\textnormal{{adoptCommit}}_{2},\dots$, where $\textnormal{{adoptCommit}}_{r}$ tests for agreement in round $r$. The input to the consensus procedure is the usual consensus input input, plus the hidden-threshold function $f$ and security parameter $s$ for thresholdConciliator plus a nonce value used to prevent replay attacks on oracleConciliator. The oracleConciliator procedure assumes access to a random oracle that takes as input the current round number and a nonce value, and returns 00 or $1$ with equal probability.
Theorem 7.1.
Let $t$ be the cost of computing $f$ and $t^{\prime}$ the cost of computing $H$. Then Algorithm 7.1 satisfies agreement and validity in all executions, and there is a constant $c>0$ such that if the adversary is limited to $cst$ time complexity, Algorithm 7.1 terminates in expected $O\left(s+2^{-s/2}st/t^{\prime}\right)$ steps and $O\left(t+2^{-s/2}st\right)$. time for each process.
If the time complexity of the adversary is bounded by a larger value $b$, then Algorithm 7.1 still terminates in expected $O(b/t^{\prime})$ steps and $O(b)$ time for each process.
Proof.
Agreement and validity follow immediately from the properties of the adopt-commit objects and the fact that each conciliator satisfies validity. (For oracleConciliator, this is because if all processes have the same input, all process see $\textnormal{{present}}[r][¬v]=0$ in line 7.1 and return this input.)
If the adversary is limited to at most $cst$ time complexity, then from Corollary 5.3, the algorithm finishes after the first round with probability at least $1-2^{-s/2}$, incurring a cost of $O(s)$ steps and $O(t)$ time to each process. In the event that it does not finish, the adversary loses its ability to compute $H(r,\textnormal{{nonce}})$ after $O(cst/t^{\prime})$ rounds, after which each round is the last with probability $1/2$, since there is at most one value $v$ for which the processes with value $v$ skip the oracle, and the oracle output equals $v$ with this probability. It follows that in this case, the processes consume an additional $O(cst/t^{\prime})$ steps and $O(cst)$ time. Multiplying by the bound on the probability of this case gives the expected step complexity for each process of $O\left(s+2^{-s/2}st/t^{\prime}\right)$ and expected time complexity of $O\left(t+2^{-s/2}st\right)$.
If the adversary is given a sufficiently large time bound $b$, it can instead allocate $O(st)$ time to force the round-$1$ conciliator to fail. Predicting the output of $H$ in each subsequent round requires $t^{\prime}$ time, equal up to small additive constant to the time cost to each process that executes $H$. But after $O(b/t^{\prime})$ rounds, the adversary again runs out of time, and so the protocol terminates with an expected cost of $O(b/t^{\prime})$ steps and $O(b)$ time for each process. ∎
If the algorithm designer can tune $s$ based on a known bound $b$ on the time complexity of the adversary, the natural choice is to choose $s$ and $t$ so that the first case of the theorem holds adversary pays $Θ(st)$ time (assuming $2^{-s/2}s=O(1)$) while each process pays only $O(t)$.
8 Possible implementations of random oracles
Let nonce be a (tunable) part of $P[1,\ldots,\ell]$, for a given round $r$ consider $F=SHA(r,\textnormal{{nonce}})$, where the number of bits in $F$ is significantly smaller than the number of bits of the chosen nonce, allowing (by the pigeon hole principle) a set of preimages $PI$ (that includes $r,\textnormal{{nonce}}$) to yield the same value of $F$. Define $H(r,\textnormal{{nonce}})$ to be the parity of the (lexicographically) smallest preimage in $PI$.
The length of the portion used from, say, $P[1,\ldots,\ell]$ (the $C$ array, or the common program code, can serve us too) may tune the exhaustive search for the smallest preimage, as an exhaustive search over longer potential preimages may require more computation resources. Another obvious computation tuning capability is the number of output bits of the cryptographic hash function, e.g., SHA128 where the output is 128 bits, or one that outputs a different number of bits.
Processes use this computationally demanding procedure to harvest (pseudo) random bits, forcing the scheduler to invest more computation to coordinate a bivalent preserving strategy.
9 Concluding remarks
@@
Acknowledgment: It is a pleasure to thank Sergio Rajsbaum for the initial interactions.

### v2

@@ 6.3.1 Outline
2 $C[1,...,\ell][1,...m-1]$ an uninitizalized array of bit strings
@@ 6    $P[i]\leftarrow$ random bit string of length $k$
7    $r_{i}\leftarrow$ random bit string of length mkm-k
8    $d\leftarrow H_{m}(P[i]\|r_{i}$)
9    $C[i]\leftarrow d_{1}d_{2}\dots d_{m-1}$
10    $V[i]\leftarrow d_{m}\oplus f_{T}(i)$
11 return $C,P,V$
12
Algorithm 6.1 Threshold Encapsulation
6.3.2 Preprocessing Algorithm
Algorithm 6.1 is executed by the programmer to obtain an obfuscated protocol. The integer parameters $\ell$, $T$, $k$, and $m$ are declared (Line 1). Where $\ell$ is the span in which a threshold $t$ is defined, $m$ is the number of bits used by the one-way function, say 256 for SHA-256, and $k$ is the length of the (exposed) prefix nonce of the $m$ bits input (and output) of the SHA.
The for loop (lines 6.1 to 6.1) iterates over all possible values of $i$ and computes the cryptographic hash (e.g., the common one-way function such as SHA) over the nonce concatenated with $r_{i}$, a concealed random string of length mkm-k. The computation starts (line 6.1) with a selection of random bit string of size $k$ as the nonce for $i$ and entry and stores it in $P[i]$. Next, the algorithm randomly selects a bit string $r_{i}$ (line 6.1); Then, the algorithm computes the hash with the concatenation of the two strings as input (line 6.1). Then stores the first m1m-1 bits of the hash in $C[i]$ (line 6.1), and stores the XOR of the last bit with $f_{T}(i)$ in $V[i]$ (line 6.1). By the time the algorithm reaches the return command (line 6.1), it has computed the hashes for all the $\ell$ hash values of $f_{T}$, and all the nonces used to calculate them are stored in $P$; thus, the algorithm returns them to be embedded in the protocol program.
local data: integers $\ell,m,k$, precomputed array of hashes $C[1,...,\ell][1,...,m-1]$, prefixes $P[1,...,\ell][1,...,k]$, and bits $V[1,\dots,\ell]$
1 procedure $\textnormal{{probe}}(i\in[\ell])$
2 Y[1,,mk]Y[1,...,m-k] an uninitialized bit string
3 D[1,,m]D[1,...,m] an uninitialized bit string
4 $Y\leftarrow$ string that satisfies $C[i][1,\dots,m-1]=H_{m}(P[i]\|Y)$
5 $D\leftarrow H_{m}(P[i]\|Y)$
6 return $D[m]\oplus V[i]$
@@ 6.3.2 Preprocessing Algorithm
6.3.3 Probing algorithm
The processes use the probe procedure described in Algorithm 6.2 to compute $f_{T}(i)$.
@@ 6.3.3 Probing algorithm
The algorithm works by finding the pre-image $P[i]\|r_{i}$ that was used to compute $C[i]$ in the preprocessing phase. It begins by initializing a bit string $Y$ of length mkm-k to hold potential suffixes to pre-image $P[i]$ of $C[i]$, and an $m$-length bit string $D$ to hold the hash that is equal to $C[i]$ in the first m1m-1 bits. Next, it finds and stores in $Y$ a suffix for the pre-image of $C[i]$ such that the first m1m-1 bits of the hash of $P[i]\|Y$ are equal to $C[i]$ (line 6.2). It stores this hash in $D$ (line 6.2). This means that, with high probability $D[m]=H_{m}(P[i]\|r_{i})[m]$. The exact method of finding this pre-image has not been specified, but exhaustive search mining or any other method to find a pre-image of a cryptographic hash function can be employed. Thus, Algorithm 6.2 returns $f_{T}(i)$ by computing $D[m]\oplus V[i]$ (line 6.2), since the preprocessing algorithm defined $V[i]=H_{m}(P[i]\|r_{i})[m]\oplus f_{T}(i)$, the algorithm returns $f_{T}(i)$.
7 Conclusion